Transverse waves with a speed of 44.0 m/s are to be produced on a stretched string. A 5.10-m length of string with a total mass of 0.0600 kg is used. Find (a) What is the required tension in the string? InN

(b) Calculate the wave speed in the string if the tension is 8.00 N.
m/s
Answer #1

a) Wave speed in stretched string   

Here v=44.0 m/s , =M/L =0.0600/5.10 =0.01176 kg/m

So

T =v2 * =442 (0.01176) =22.76736 N

T=22.77 N Answer

b)

Answer